ICM 1998, Berlin, Aug. 18–27

Abstracts of Short Communications and Posters

Section:

8. Analysis

1991 MS Classification: 30C45 Owa, Shigeyoshi, Department of Mathematics, Kinki University, Japan Some Properties for Convulutions Let A denote the class of analytic functions f(z) normalized by f(0) = f′(0)−l = 0 in the open unit disk. The object of the present talk is to show some properties of convolutions (or Hadamard products) of functions in A.